OpenStudy (anonymous):

find tangent to witch of agnesi y= 8/(4+x^2) at point (2,1)

OpenStudy (anonymous):

did you find the derivative?

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{1}{f(x)}=\frac{-f'(x)}{f^2(x)}\] is a start

OpenStudy (anonymous):

would the derivative be y= (-16x)/(x^2+4)^2

OpenStudy (anonymous):

yes, looks good

OpenStudy (anonymous):

how to i proceed?

OpenStudy (anonymous):

find the slope by taking \(f'(2)\) then use the point slope formula for the equation of the line

OpenStudy (anonymous):

so would it be y=1/2x

OpenStudy (anonymous):

i don't know i didn't do it, but i can check

OpenStudy (anonymous):

slope should be negative though

OpenStudy (anonymous):

i get a slope of \(-\frac{1}{2}\) and the line as \(y=-\frac{1}{2}x+2\)

OpenStudy (anonymous):

how do i make the tangent of that line at (2,1)???

OpenStudy (anonymous):

i used \(y-y_1=m(x-x_1)\) with \(m=-\frac{1}{2},x_1=2,y_1=1\)

OpenStudy (anonymous):

when i graphed it though, the tangent was at a point 2, 7 or something strange like that